What point on x-axis is equidistant from (7,  6) and (-3, 4).
 
 

Grade:11

1 Answers

Aman kumar
5 years ago
Let  P(x,0) be a point on x-axis which is equidistant from the given points.
So from distance formulae,
[(x-7)^2+(0-6)^2]^1/2=[(-3-x)^2+(4-0)^2]^1/2
(x-7)^2+(0-6)^2=(-3-x)^2+(4-0)^2
x^2+49-14x+36=9+x^2+6x+16
49+36-9-16=6x+14x
20x=60
x=3
So, the required point is (3,0)